The first term of an geometric progression is 12 and the fifth term is 55. Determine the 8th term and the 11th term.
Answer by mananth(16946)

The first term of an geometric progression is 12 and the fifth term is 55. Determine the 8th term and the 11th term.
a1 =12
an=ar^(n-1)
a5 = 12*r(^4)
55 =12*r^(4)
r^4=55/12
r=1.4631....
a=12, r = 1.4361
Use formula
a8 =12*(1.7631)^7